OpenStudy (anonymous):

what are the 5 terms for this sequence?? : f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. please help! thanks! <33

myininaya (myininaya):

The first 5 terms? Which terms are you looking for?

myininaya (myininaya):

Like if we plug in 3, what do we get?

myininaya (myininaya):

Replace n with 3

myininaya (myininaya):

f(n) = f(n - 1) + f(n - 2) f(3)=f(3-1)+f(3-2) f(3)=f(2)+f(1) You know that f(2)=1 and f(1)=0 So f(3)=?

OpenStudy (anonymous):

Umm.. okay. Let's see. F(3) equals... 1?

myininaya (myininaya):

Yep now what is f(4)=?

myininaya (myininaya):

f(n) = f(n - 1) + f(n - 2) replace n with 4 like so: f(4)=f(4-1)+f(4-2) f(4)=f(3)+f(2)

myininaya (myininaya):

So f(4)=?

OpenStudy (anonymous):

f(4) equals 2?

myininaya (myininaya):

yep And you can do f(5) now and so on...

OpenStudy (anonymous):

Ohh. Okay! Can you help me understand another one? You describe it much simpler than the lesson does. :))

myininaya (myininaya):

Ok, go for it!

OpenStudy (anonymous):

(i got the first question mixed up but //this// is the one that needs 5 terms in the sequence! woops! :o) f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (anonymous):

oh damn i did it again. okay i meant: If g(x) = 2(x - 4), find the value of x if g(x) = 20. ugh! LOL woopsie (again)

myininaya (myininaya):

So g(x)=2(x-4) But you want to find what x is if g(x)=20 So 20=g(x)=2(x-4) Since all of those these should be the same then 20 has the to be the same as 2(x-4) So we have 20=2(x-4) solve for x.

OpenStudy (anonymous):

Uh... 14?

myininaya (myininaya):

yep gj